### v1

# On a question of Blecher, Pisier, Shlyakhtenko

Abstract.
We show the failure of a matricial version of Grothendieck’s theorem for operator spaces, thereby resolving a long-standing open question in the field. Moreover, by showing that such a counterexample can occur in the simplest context of commutative $C^{*}$-algebras, we address some other open questions in operator algebras. Our constructions, completely explicit and fairly simple, are inspired by some techniques in quantum information theory.
Key words and phrases:
Grothendieck’s Theorem, Operator spaces
MJ is partially supported by the NSF grants DMS 2247114 and Raise-TAG183917. C.P. is partially supported by the MICINN project PID2020-113523GB-I00, by QUITEMAD+-CM, P2018/TCS4342, funded by Comunidad de Madrid and by Grant CEX2019-000904-S funded by MCINN/AEI/ 10.13039/501100011033. C.P. gratefully acknowledges financial support for this publication by the Fulbright Program, which is sponsored by the U.S. Department of State, the U.S.- Spain Fulbright Commission.
1. Introduction and main results
Grothendieck’s contributions to functional analysis [8], despite initially overlooked, transformed Banach space theory and gave rise to the Grothendieck program on operator algebras [18]. Moreover, Grothendieck’s “fundamental theorem” has applications today that span a wide range of fields, including computer science and quantum information theory [13, 22].
In his work, Grothendieck employed a functorial approach to study the smallest $\otimes_{\epsilon}$ and largest $\otimes_{\pi}$ norms on the tensor product of two Banach spaces. His fundamental theorem can be formulated as follows:
(1.1) $\ell_{1}\otimes_{\epsilon}\ell_{1}\cong\ell_{1}\otimes_{\gamma_{2}^{*}}\ell_{1}\quad\mbox{isomorphically}\,,$
where $\gamma_{2}^{*}$ is dual of the $\gamma_{2}$ norm, derived from the Hilbert space factorization norm for linear maps. This theorem highlights a remarkable property of tensor products of Banach spaces, revealing deep connections between different norms and structures (see [5] and the references therein).
In 1974 Pisier provided a positive answer to one of Grothendieck’s problems by showing that
(1.2) $A^{*}\otimes_{\epsilon}B^{*}\cong A^{*}\otimes_{\gamma_{2}^{*}}B^{*}\quad\mbox{isomorphically}$
holds for $C^{*}$-algebra $A$, $B$ with an additional approximability assumption [16]. That is, Pisier replaced commutative $L_{1}$-spaces by noncommutative ones. This result was later improved by Haagerup [9], who removed the approximability assumption made in [16] and clarified the correct analogue to $\gamma_{2}^{*}$ in the context of operator algebras.
The category of operator spaces, which are closed subspaces of $C^{*}$-algebras endowed with an induced matricial structure, is closed under taking dual spaces, closed subspaces and quotients. In particular, the dual of a $C^{*}$-algebra has a natural operator space structure. Moreover, a theory of tensor products can be naturally developed in this category [3], allowing for analogues of Grothendieck’s smallest and largest tensor norms. It turns out that for operator spaces $X\subset\mathcal{B}({H})$ and $Y\subset\mathcal{B}(K)$, the smallest norm
$X\otimes_{\min}Y\subset\mathcal{B}(H\otimes_{2}K)$
matches exactly the construction for tensor products of $C^{*}$-algebras. On the other hand, Grothendieck’s Hilbert space factorization norm is usually replaced by the so-called Haagerup tensor product
$\|\xi\|_{X\otimes_{h}Y}\,=\,\inf_{\xi=\sum_{k=1}^{m}x_{k}\otimes y_{k}}\left\|(x_{1},\cdots,x_{m})\right\|\left\|\left(\begin{array}[]{c}y_{1}\\ \vdots\\ y_{m}\end{array}\right)\right\|\,.$
The Haagerup tensor product $X\otimes_{h}Y$ comes with a natural operator space structure and appears to be the appropriate tool to refine (1.2) into a form more closely resembling (1.1). Indeed, in 1991 Blecher conjectured a fully functorial operator space version of Grothendieck’s theorem: for general $C^{*}$-algebras $A$ and $B$, it holds
(1.3) $A^{*}\otimes_{\min}B^{*}\,\stackrel{{\scriptstyle?}}{{\cong}}A^{*}\otimes_{\mu}B^{*}\quad\mbox{completely isomorphically}\,,$
where $A^{*}\otimes_{\mu}B^{*}=A^{*}\otimes_{h}B^{*}+B^{*}\otimes_{h}A^{*}$ may be considered as a quotient of the direct sum $A^{*}\otimes_{h}B^{*}\oplus B^{*}\otimes_{h}A^{*}$ with respect to the map $q(a_{1}\otimes b_{1}\oplus b_{2}\otimes a_{2})=a_{1}\otimes b_{1}+a_{2}\otimes b_{2}$.
A breakthrough in this problem was achieved by Piser and Shlyakhtenko in [20], where they conclusively demonstrated that the answer to the preceding question is indeed positive at the Banach space level. They proved
(1.4) $A^{*}\otimes_{\min}B^{*}\cong A^{*}\otimes_{\mu}B^{*}\quad\mbox{isomorphically}\,,$
assuming that the $C^{*}$-algebras $A$ and $B$ are exact. Notably, their results extended beyond the setting of $C^{*}$-algebras to encompass general exact operator spaces. Subsequently, Haagerup and Musat in [10] furthered this understanding to include general $C^{*}$-algebras. However, the broader question concerning the operator space structure of the corresponding tensors products in Conjecture (1.3) remained unresolved in [20] (see also [18, Problem 21.2]).
The first result presented in this work provides a negative answer to Blecher’s conjecture. Moreover, to this end, we do not consider the whole matricial structure of the tensor product, but only a column structure. If $\mathcal{K}(\ell_{2})$ denotes the space of compact operators on $\ell_{2}$ and $C$ is the column operator space, we prove:
Theorem 1.1.
$C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\min}\mathcal{K}(\ell_{2})^{*})\neq C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\mu}\mathcal{K}(\ell_{2})^{*})\quad\text{isomorphically}\,,$
In fact, we will demonstrate an even stronger result than Theorem 1.1 by proving the same statement with the operator space $\mathcal{K}(\ell_{2})^{*}\otimes_{\min}\mathcal{K}(\ell_{2})^{*}$ replaced by $\mathcal{K}(\ell_{2})^{*}\otimes_{\max}\mathcal{K}(\ell_{2})^{*}$. This allows us to link the previous conjecture on Grothendieck’s theorem with some fundamental questions in operator algebras. To explain the scope of this and the following results, given an operator space $X$, let us denote by $OA(X)$ its unital universal operator algebra and by $C^{*}\langle X\rangle$ its unital universal $C^{*}$-algebra, respectively (see Section 2 for details). Then, it was proved in [14] that
$X\otimes_{\mu}Y\subset OA(X)\otimes_{max}OA(Y)\quad\mbox{completely isometrically},$
while one defines the max norm so that
$X\otimes_{max}Y\subset C^{*}\langle X\rangle\otimes_{max}C^{*}\langle Y\rangle\quad\mbox{completely isometrically}.$
Hence, the analogous statement to Theorem 1.1, when the min is replaced with the max norm, can be restated in terms of the corresponding subspaces of the previous algebras when $X=Y=\mathcal{K}(\ell_{2})^{*}$ is the space of trace class operators.
The previous viewpoint strongly motivates considering Conjecture 1.3 when focusing on commutative $C^{*}$-algebras; specifically, when $X=Y=\ell_{1}$. This problem, already considered in [20] (see also [17, Chapter 25]), highlights the equivalence, at the matrix level, to the context originally considered by Grothendieck, since the equivalence
$\ell_{1}\otimes_{\mu}\ell_{1}\cong\ell_{1}\otimes_{max}\ell_{1}\quad\mbox{isomorphically}\,,$
follows from the classical Grothendieck’s Theorem stated in (1.1). Our second result, the main result of this work, states that this is not longer true in the category of operator spaces:
Theorem 1.2.
$\mathcal{K}(\ell_{2})\otimes_{\min}(\ell_{1}\otimes_{max}\ell_{1})\neq\mathcal{K}(\ell_{2})\otimes_{\min}(\ell_{1}\otimes_{\mu}\ell_{1})\quad\text{isomorphically}\,.$
Since it is well known that $C^{*}\langle\ell_{1}\rangle=C^{*}(\mathbb{F}_{\infty})$, one can alternatively express Theorem 1.2 by indicating that the natural inclusion $OA(\ell_{1})\otimes_{max}OA(\ell_{1})\subseteq C^{*}(\mathbb{F}_{\infty})\otimes_{max}C^{*}(\mathbb{F}_{\infty})$ is not completely isomorphic, even when restricted to the linear subspace $\ell_{1}\otimes\ell_{1}$. According to the precise definitions of $OA(\ell_{1})$ and $C^{*}\langle\ell_{1}\rangle$, Theorem 1.2 underscores a significant limitation in lifting specific sets of commuting contractions to commuting unitaries (see Remark 4.1).
The approach followed in this work delves into the deep relationship between the theories of operator spaces and nonlocal games [15]. On one hand, the proof of Theorem 1.1 is inspired by the CHSHn games [1], an extension of the renowned CHSH inequality, which holds paramount significance in quantum theory. On the other hand, the proof of Theorem 1.2 utilizes certain embeddings between noncommutative $L_{p}$-spaces proved by the second and third authors in [12], which are motivated by the quantum teleportation protocol in quantum computation. The combination of these techniques, along with some operator space computations, lead to completely explicit and (arguably) simple elements to demonstrate the previous theorems.
2. preliminaries
2.1. Basic on operator spaces and operator algebras
In this section we will review some of the necessary basics of operator spaces and their tensor products which we will invoke throughout the manuscript. We refer to [6, 17] for comprehensive references on the topic.
An operator space is a closed subspace $X\subset\mathcal{B}(H)$, where $\mathcal{B}(H)$ denotes the spaces of bounded operators acting on a complex Hilbert space $H$. For any such subspace one naturally obtains a sequence of matrix norms $\|\cdot\|_{n}$ on $M_{n}(X)$ via the inclusion $M_{n}(X)\subseteq M_{n}(\mathcal{B}(H))\simeq\mathcal{B}(H^{\oplus n})$. Ruan’s Theorem [21] characterizes the sequences of norms that can be achieved in this manner, offering an alternative definition of an operator space as a complex Banach space $X$ equipped with a sequence of matrix norms $(M_{n}(X),\|\cdot\|_{n})$ that satisfy specific conditions. In both scenarios, we say that $X$ is endowed with an operator space structure.
When considering operator spaces, the norm on linear operators must reflect the matrix structure defined by the operator space structure. For a linear map between operator spaces $T:X\rightarrow Y$, we say that it is completely bounded if $\|T\|_{cb}:=\sup_{n}\|\text{Id}\otimes T:M_{n}(X)\rightarrow M_{n}(Y)\|$ is finite. We will say that $T$ is completely contractive if $\|T\|_{cb}\leq 1$. Moreover, $T$ is termed a complete isomorphism (resp. complete isometry) if it is an isomorphism with $T$ and $T^{-1}$ completely bounded (resp. completely contractive).
Given an operator space $X$, then $X^{*}$ has a privileged operator space structure with norms on $M_{n}(X^{*})$ given via the identification
$M_{n}(X^{*})=\text{CB}(X,M_{n}),$
where the latter denotes the operator space of completely bounded maps $T:X\to M_{n}$ endowed with the norm $\|\cdot\|_{cb}$. Every $C^{*}$-algebra $A$ has a natural operator space structure. Therefore, by duality, $\ell_{1}$ and the space of trace class operators $S_{1}$ also have a natural operator space structure, as they are the duals of $c_{0}$ and the space of compact operators $\mathcal{K}(\ell_{2})$, respectively. Moreover, the natural operator spaces structure on the space of bounded operators acting on $\ell_{2}$, $S_{\infty}$, matches the one given by the duality $S_{1}^{*}=S_{\infty}$.
Important examples which we use throughout this manuscript are the column and row Hilbert operator spaces. Given a Hilbert space $H$, the column operator space $H_{c}$ is defined by means of the natural identification
$H\cong\mathcal{B}({\mathbb{C}},H).$
For the particular case $H=\ell_{2}^{n}$, we will simply denote $C_{n}$. By duality, we can define the row operator space $H_{r}=(H_{c})^{*}$, where we recall that, given a Hilbert space $H$, its dual space $H^{*}$ is naturally identified with the conjugate $\overline{H}$. As in the previous case, the finite dimensional row space will be simply denoted by $R_{n}$.
An abstract theory for operator space tensor products was first developed by Blecher and Paulsen in [3] in which they introduce the projective and injective operator space tensor products. Given two operator spaces $X,Y$, and $w\in M_{n}(X\otimes Y)$, define the operator space projective norm of $w$ as
$\|w\|_{\wedge}:=\inf\{\|\alpha\|\|x\|\|y\|\|\beta\|\},$
where the infimum is taken over $p,q,n\in\mathbb{N}$, $x\in M_{p}(X)$, $y\in M_{q}(Y)$, $\alpha\in M_{n,pq}$, $\beta\in M_{pq,n}$, such that $w=\alpha(x\otimes y)\beta$. We denote by $X\hat{\otimes}Y$ the corresponding operator space. As one may expect from its name, the projective tensor norm preserves (complete) quotients. One may also consider this operator space as the “predual” of the operator space $\text{CB}(X,Y^{*})$. In particular, one has the complete isometry $(X\hat{\otimes}Y)^{*}=\text{CB}(X,Y^{*})$. As previously stated, the operator space injective (minimal) tensor product is induced via the embedding into the minimal C*-algebra tensor product. Thus, if $X\subset\mathcal{B}(H),Y\subset\mathcal{B}(K)$, then we have the completely isometric embedding $X\otimes_{min}Y\subset\mathcal{B}(H\otimes_{2}K)$. In particular, it follows that the min norm is preserved by (complete) isometries. It can be seen that, given $w\in M_{n}(X\otimes Y)$, one has
$\|w\|_{M_{n}(X\otimes_{min}Y)}=\sup\Big{\|}\big{(}\text{Id}\otimes T\otimes S\big{)}(w)\Big{\|}_{M_{n}(\mathcal{B}(H\otimes K))},$
where the supremum is taken over all Hilbert spaces $H$ and $K$ and all complete contractions $T:X\rightarrow\mathcal{B}(H)$ and $S:Y\rightarrow\mathcal{B}(K)$. In comparison to the projective tensor product, we have the completely isometric embedding $X\otimes_{min}Y\subseteq CB(X^{*},Y).$ Hence, we see that the projective and the injective norms are dual to each other and for finite-dimensional operator spaces $X$, $Y$, there are the following completely isometric identifications
$(X\hat{\otimes}Y)^{*}=X^{*}\otimes_{min}Y^{*},\hskip 11.38092pt(X\otimes_{min}Y)^{*}=X^{*}\hat{\otimes}Y^{*}.$
Given operator spaces $X\subset\mathcal{B}(H)$ and $Y\subset\mathcal{B}(K)$, we define the Haagerup operator space norm for $w\in M_{n}(X\otimes Y)$ as
$\|w\|_{h}:=\inf\{\|u\|_{M_{n,r}(X)}\|v\|_{M_{r,n}(Y)}\},$
where the infimum runs over $r\in\mathbb{N}$ and $u$, $v$, such that $w_{i,j}=\sum_{k=1}^{r}u_{i,k}\otimes v_{k,j}$ for every $i,j$. We denote by $X\otimes_{h}Y$ the corresponding operator space, which is usually called the Haagerup operator space tensor product. Despite being injective, projective and self-dual, it has been long known that $\otimes_{h}$ is not commutative.
It is well known that the previous norms behave nicely when considering tensor products of linear maps. Specifically, if $\alpha$ denotes any of the previous three norms and $T_{i}:X_{i}\rightarrow Y_{i}$ are completely bounded maps for $i=1,2$, then the map
$T_{1}\otimes T_{2}:X_{1}\otimes_{\alpha}X_{2}\rightarrow Y_{1}\otimes_{\alpha}Y_{2}$
is completely bounded and
$\|T_{1}\otimes T_{2}\|_{cb}=\|T_{1}\|_{cb}\|T_{2}\|_{cb}.$
This is usually refer to as the metric mapping property in the category of operator spaces. Also, by iterating the process, one can define the aforementioned norms on the tensor product of more than two spaces. Remarkably, all three norms exhibit associativity. Furthermore, the previous properties (related to projectivity, injectivity, duality, etc.) remain consistent and apply in the same manner.
The previously discussed tensor norms are closely related when examining the column and row spaces. In fact, it can be proven that for every operator space $X$, the equalities
[…]
hold completely isometrically. By duality, one also has $R\hat{\otimes}X=R\otimes_{h}X$ and $X\hat{\otimes}C=X\otimes_{h}C$ completely isometrically.
The previous identifications, along with the aforementioned properties of the tensor norms we are considering, easily imply that the following identities are complete isometries:
[…]
Equation (2.2) and the well-known estimate $\|\text{Id}:R_{n}\rightarrow C_{n}\|_{cb}=\|\text{Id}:C_{n}\rightarrow R_{n}\|_{cb}=\sqrt{n}$, can be used to show that for every operator space $X$, it holds:
(2.3) $\Big{\|}\frac{\text{Id}}{n}:S_{\infty}^{n}\otimes_{min}X\rightarrow S_{1}^{n}\hat{\otimes}X\Big{\|}_{cb}=1.$
As emphasized in the introduction, the non-commutativity of the Haagerup tensor norm leads one to consider the so-called symmetrized Haagerup tensor product $\otimes_{\mu}$, where given $w\in M_{n}(X\otimes Y)$ we define
(2.4) $\|w\|_{\mu}:=\inf\{\|w_{1}\|_{M_{n}(X\otimes_{h}Y)}+\|w_{2}^{T}\|_{M_{n}(Y\otimes_{h}X)}:w=w_{1}+w_{2}\}.$
Here, for $w=\sum_{i}a_{i}\otimes x_{i}\otimes y_{i}$, we have let $w_{2}^{T}=\sum_{i}a_{i}\otimes y_{i}\otimes x_{i}$.
Given an operator space $X$, its universal (unital) operator algebras $OA(X)$ (see [17, Chapter 6] for a detailed construction of the algebra) is characterized by the following universal property: for any complete contraction $T:X\rightarrow\mathcal{B}(H)$ there exists a unique unital morphism $\pi:OA(X)\rightarrow\mathcal{B}(H)$ extending $T$, where $X$ is naturally embedded in $OA(X)$. This algebra is particularly interesting since, for operator spaces $X$ and $Y$, it is known [14, Lemma 5] that
$X\otimes_{\mu}Y\subset OA(X)\otimes_{max}OA(Y)\hskip 8.5359pt\text{completely isometrically.}$.
Here, the max norm is defined at the level of unital operator algebras. In particular, the previous inclusion shows that for any element $w\in M_{n}(X\otimes Y)$, we have
$\|w\|_{M_{n}(X\otimes_{\mu}Y)}=\sup\Big{\|}\big{(}\text{Id}\otimes T\odot S\big{)}(w)\Big{\|}_{M_{n}(\mathcal{B}(H))},$
where the supremum runs over all complex Hilbert spaces $H$ and all complete contractions $T:X\rightarrow\mathcal{B}(H)$ and $S:Y\rightarrow\mathcal{B}(H)$ with commuting ranges and where we have denoted $T\odot S$ the operator such that $T\odot S(x\otimes y)=T(x)S(y)$.
On the other hand, the universal (unital) $C^{*}$-algebra associated to an operator space $X$, $C^{*}\langle X\rangle$ (see [17, Chapter 8] for a detailed construction of the algebra [^1] This algebra should not be confused with the injective envelop of an operator space $I(X)$ as discussed in [4, 11].), is characterized by the following universal property: For any complete contraction $T:X\rightarrow\mathcal{B}(H)$ there exists a unique unital *-homomorphism $\pi:C^{*}\langle X\rangle\rightarrow\mathcal{B}(H)$ extending $T$, where $X$ is naturally embedded in $C^{*}\langle X\rangle$. It can be seen that $OA(X)$ can be identified with the closed subalgebra generated by $X$ in $C^{*}\langle X\rangle$ so that $OA(X)\subset C^{*}\langle X\rangle$ completely isometrically. One then defines the max tensor product of two operator spaces $X$ and $Y$ so that the following inclusion is a complete isometry:
$X\otimes_{max}Y\subset C^{*}\langle X\rangle\otimes_{max}C^{*}\langle Y\rangle,$
where $C^{*}\langle X\rangle\otimes_{max}C^{*}\langle Y\rangle$ denotes the corresponding tensor product in the category of $C^{*}$-algebras. In particular, if we use bracket notation $[A,B]$ for the commutator of two operators in $\mathcal{B}(H)$, for any element $w\in M_{n}(X\otimes Y)$ we have
(2.5) $\|w\|_{M_{n}(X\otimes_{max}Y)}=\sup\Big{\|}\big{(}\text{Id}\otimes T\odot S\big{)}(w)\Big{\|}_{M_{n}(\mathcal{B}(H))},$
where the supremum is taken over all complex Hilbert spaces $H$ and all complete contractions $T:X\rightarrow\mathcal{B}(H)$ and $S:Y\rightarrow\mathcal{B}(H)$ such that $[T(x),S(y)]=[T(x),S(y)^{\dagger}]=0$ for every $x\in X$ and $y\in Y$. While the universal algebra is in general an unknown object, it is well known that $C^{*}\langle\ell_{1}^{n}\rangle=C^{*}(\mathbb{F}_{n-1})$ and $C^{*}\langle S_{1}^{n}\rangle=B_{n}$, where $C^{*}(\mathbb{F}_{n-1})$ is the full $C^{*}$-algebra associated to the free group and $B_{n}$ denotes the Brown algebra.
2.2. Some basic results involving the Haagerup tensor norm
In this section, we present some auxiliary results that will significantly simplify the readability of the proofs for the main theorems.
Lemma 2.1.
Given any natural number $n$, we have
$\|\text{Id}:S_{1}^{n}\hat{\otimes}(S_{1}^{n}\otimes_{min}S_{\infty}^{n})\rightarrow S_{1}^{{n^{2}}}\otimes_{min}S_{\infty}^{n}\|_{cb}\leq 1.$
Proof.
According to the properties of the min and the Haagerup tensor norms, the following identifications are complete isometries.
$S_{1}^{n}\hat{\otimes}(S_{1}^{n}\otimes_{min}S_{\infty}^{n})=S_{1}^{n}\hat{\otimes}(S_{\infty}^{n}\otimes_{min}S_{1}^{n})$ $=R_{n}^{A}\otimes_{h}(C_{n}^{B}\otimes_{h}(R_{n}^{C}\otimes_{h}C_{n}^{C})\otimes_{h}R_{n}^{B})\otimes_{h}C_{n}^{A}$
$=(R_{n}^{A}\otimes_{h}C_{n}^{B})\otimes_{h}R_{n}^{C}\otimes_{h}C_{n}^{C}\otimes_{h}(R_{n}^{B}\otimes_{h}C_{n}^{A}).$
Note that we have used the super indexes $A$, $B$ y $C$ to denote the three spaces appearing in the tensor product $S_{1}^{n}\hat{\otimes}(S_{\infty}^{n}\otimes_{min}S_{1}^{n})$. This allows us to indicate to which of these spaces the column or row spaces correspond in the following equalities.
Now, since $\|\text{Id}:R_{n}^{A}\otimes_{h}C_{n}^{B}\rightarrow C_{n}^{B}\otimes_{h}R_{n}^{A}\|_{cb}\leq 1$ and $\|\text{Id}:R_{n}^{B}\otimes_{h}C_{n}^{A}\rightarrow C_{n}^{A}\otimes_{h}R_{n}^{B}\|_{cb}\leq 1$, we conclude our proof by noticing the following completely isometric identifications:
$(C_{n}^{B}\otimes_{h}R_{n}^{A})\otimes_{h}R_{n}^{C}\otimes_{h}C_{n}^{C}\otimes_{h}(C_{n}^{A}\otimes_{h}R_{n}^{B})$ $=C_{n}^{B}\otimes_{h}(R_{n}^{A}\otimes_{h}R_{n}^{C}\otimes_{h}C_{n}^{C}\otimes_{h}C_{n}^{A})\otimes_{h}R_{n}^{B}$
$=C_{n}^{B}\otimes_{h}(R_{{n^{2}}}^{AC}\otimes_{h}C_{{n^{2}}}^{AC})\otimes_{h}R_{n}^{B}$
$=S_{\infty}^{n}(S_{1}^{{n^{2}}}).$
Lemma 2.2.
Given two operator spaces $X$ and $Y$ and any natural number $n$, we have
  1. a)
    $\|T:X\otimes_{h}M_{n}(Y)\rightarrow M_{n}(X\otimes_{h}Y)\|_{cb}\leq 1$, where $T$ is defined as $T(x\otimes A\otimes y)=A\otimes x\otimes y$ on elementary tensors.
  2. b)
    $\|\text{Id}:M_{n}(X)\otimes_{h}Y\rightarrow M_{n}(X\otimes_{h}Y)\|_{cb}\leq 1.$
Proof.
In order to prove the first estimate, note that
$X\otimes_{h}M_{n}(Y)=(X\otimes_{h}(C_{n}\otimes_{h}Y\otimes_{h}R_{n}))=(X\otimes_{h}C_{n})\otimes_{h}(Y\otimes_{h}R_{n})$
completely isometrically. Now, $\|\text{Id}:X\otimes_{h}C_{n}\rightarrow C_{n}\otimes_{h}X\|_{cb}\leq 1$. On the other hand,
$(C_{n}\otimes_{h}X)\otimes_{h}(Y\otimes_{h}R_{n})=C_{n}\otimes_{h}(X\otimes_{h}Y)\otimes_{h}R_{n}=M_{n}(X\otimes_{h}Y)$
completely isometrically. Hence, we prove the estimate a).
In order to prove the second estimate, note that
$M_{n}(X)\otimes_{h}Y=(C_{n}\otimes_{h}X\otimes_{h}R_{n})\otimes_{h}Y=(C_{n}\otimes_{h}X)\otimes_{h}(R_{n}\otimes_{h}Y)$
completely isometrically. Now, $\|\text{Id}:R_{n}\otimes_{h}Y\rightarrow Y\otimes_{h}R_{n}\|_{cb}\leq 1$. On the other hand,
$(C_{n}\otimes_{h}X)\otimes_{h}(Y\otimes_{h}R_{n})=C_{n}\otimes_{h}(X\otimes_{h}Y)\otimes_{h}R_{n}=M_{n}(X\otimes_{h}Y)$
completely isometrically. Hence, estimate b) follows. ∎
Lemma 2.3.
Let $X$ be an operator space and $(x_{k})_{k}\subset X$ any sequences. Then,
$\Big{\|}\sum_{k}e_{k}\otimes x_{k}\otimes e_{k}\Big{\|}_{R\otimes_{h}X\otimes_{h}R}=\big{(}\sum_{k}\|x_{k}\|^{2}_{X}\Big{)}^{\frac{1}{2}}.$
Proof.
Consider the operator space $Y=R\otimes_{h}X\subset\mathcal{B}(H)$, for a certain Hilbert space $H$ and denote $y_{k}=e_{k}\otimes x_{k}\in\mathcal{B}(H)$ for every $k$. It is clear that $\|y_{k}\|=\|x_{k}\|$ for every $k$. In addition, using the injectivity of the Haagerup tensor norm, we can write
$\Big{\|}\sum_{k}y_{k}\otimes e_{k}\Big{\|}^{2}_{R\otimes_{h}X\otimes_{h}R}=\Big{\|}\sum_{k}y_{k}\otimes e_{k}\Big{\|}^{2}_{\mathcal{B}(H)\otimes_{h}R}=\Big{\|}\sum_{k}y_{k}y_{k}^{*}\Big{\|}_{\mathcal{B}(H)}\leq\sum_{k}\|y_{k}\|^{2}_{\mathcal{B}(H)}=\sum_{k}\|x_{k}\|^{2}_{X},$
where we have used the triangle inequality.
In order to prove the converse inequality, note that a completely analogous proof applies if we replace $R$ with $C$ and $X$ with $X^{*}$, so that we can prove that for every sequence $(x_{k}^{*})_{k}\subset X^{*}$ the following inequality holds:
$\Big{\|}\sum_{k}e_{k}\otimes x_{k}^{*}\otimes e_{k}\Big{\|}^{2}_{C\otimes_{h}X^{*}\otimes_{h}C}\leq\sum_{k}\|x_{k}^{*}\|^{2}_{X^{*}}.$
Hence, since $(R\otimes_{h}X\otimes_{h}R)^{*}=C\otimes_{h}X^{*}\otimes_{h}C$ (completely) isometrically and also $(\ell_{2}(X))^{*}=\ell_{2}(X^{*})$ isometrically, the inequality
$\Big{(}\sum_{k}\|x_{k}\|^{2}_{X}\Big{)}^{\frac{1}{2}}\leq\Big{\|}\sum_{k}e_{k}\otimes x_{k}\otimes e_{k}\Big{\|}_{R\otimes_{h}X\otimes_{h}R},$
follows by duality. ∎
Lemma 2.4.
Let $X$ be an operator space and $(x_{k})_{k}\subset X$ any sequences. Then,
$\Big{\|}\sum_{k}e_{k}\otimes x_{k}\otimes e_{k}\Big{\|}_{R\otimes_{h}X\otimes_{h}C}=\sum_{k}\|x_{k}\|_{X}.$
Proof.
The statement follows from the isometric identification $R\otimes_{h}X\otimes_{h}C=S_{1}(X)$, simply by restricting to the diagonal of $S_{1}$, which is $\ell_{1}$. ∎
3. Proof of Theorem 1.1
In this section we prove the following theorem, from which Theorem 1.1 follows immediately.
Theorem 3.1.
For every prime number $n$ there exists an element $\eta_{n}\in C_{n^{4}}\otimes S_{1}^{n}\otimes S_{1}^{n}$ such that $\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{max}S_{1}^{n})}\leq\sqrt{2}n^{\frac{3}{4}}$ and $\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{\mu}S_{1}^{n})}\geq n$.
Here and throughout the rest of the paper, we are considering $\mathbb{Z}_{n}=\{0,1,\ldots,n-1\}$ with addition and multiplication modulo $n$. For consistency in notation, we will represent the canonical basis of ${\mathbb{C}}^{n}$ as $\{e_{i}:\,i=0,1,\ldots,n-1\}$. Additionally, we use the notation $e_{i,j}$ to denote the matrix whose entries are all zero except for a one in the $i$-th row and $j$-th column.
The upper bound in the previous result will follow from the next two lemmas.
Lemma 3.2.
Let $T:S_{1}^{n}\rightarrow\mathcal{B}(H)$ be a complete contraction and denote $E_{x}^{a}=T(e_{a,x})^{\dagger}T(e_{a,x})$ for every $x,a=1,\cdots,n$. Then, for every $x$ and every sequence of complex numbers $(\alpha_{a})_{a}$ such that $\sup_{a}|\alpha_{a}|\leq 1$, it holds that
$\Big{\|}\sum_{a}\alpha_{a}E_{x}^{a}\Big{\|}_{\mathcal{B}(H)}\leq 1.$
Proof.
According to the isometric identifications $CB(S_{1}^{n},\mathcal{B}(H))=S_{\infty}^{n}\otimes_{min}\mathcal{B}(H)=S_{\infty}(\ell_{2}^{n}\otimes_{2}{H})$, $T$ can be naturally identified with a contraction $\hat{T}=\sum_{x,a}e_{a,x}\otimes T(e_{a,x})$ so that
$\hat{T}^{\dagger}\hat{T}=\sum_{x,x^{\prime}}e_{x,x^{\prime}}\otimes\Big{(}\sum_{a}T(e_{a,x})^{\dagger}T(e_{a,x^{\prime}})\Big{)}\leq 1\!\!1_{S_{\infty}^{n}\otimes_{min}\mathcal{B}({H})}.$
In particular, it must hold that $\sum_{a}E_{x}^{a}\leq 1\!\!1_{\mathcal{B}(H)}$ for every $x$. Moreover, since the elements $E_{x}^{a}$ are semidefinite positive, the previous condition implies that, for every $x$, the linear map $T_{x}:\ell_{\infty}^{n}\rightarrow\mathcal{B}(H)$, defined as $T_{x}(e_{a})=E_{x}^{a}$ for every $a$, is a completely positive and (completely) contractive map. Thus, $\sum_{a}\alpha_{a}E_{x}^{a}=T_{x}(\sum_{a}\alpha_{a}e_{a})$ is a contraction in $\mathcal{B}(H)$. ∎
The proof of the following lemma is inspired by the analysis of the entangled value of the CHSHq games conducted in [1].
Lemma 3.3.
Let $n$ be a prime number and consider
(3.1) $\eta_{n}=\sum_{\begin{subarray}{c}x,y,a,b\in\mathbb{Z}_{n}\\ a+b=xy\end{subarray}}e_{xyab,1}\otimes e_{a,x}\otimes e_{b,y}\in C_{n^{4}}\otimes S_{1}^{n}\otimes S_{1}^{n}.$
Then,
$\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{max}S_{1}^{n})}\leq\sqrt{2}n^{\frac{3}{4}}.$
Proof.
According to Equation (2.5), we must show that for every pair of complete contractions $T:S_{1}^{n}\rightarrow\mathcal{B}({H})$ and $S:S_{1}^{n}\rightarrow\mathcal{B}({H})$ satisfying $[T(x),S(y)]=[T(x),S(y)^{\dagger}]=0$ for every $x,y\in S_{1}^{n}$, we have
$\|(\text{Id}\otimes T\odot S)(\eta_{n})\|^{2}_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{max}S_{1}^{n})}$ $\leq 2n\sqrt{n}.$
Then, for any such pair of complete contractions, it holds that
$\|(\text{Id}\otimes T\odot S)(\eta_{n})\|^{2}_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{max}S_{1}^{n})}$ $=\Big{\|}\sum_{\begin{subarray}{c}x,y,a,b\in\mathbb{Z}_{n}\\ a+b=xy\end{subarray}}e_{xyab,1}\otimes T(e_{a,x})S(e_{b,y})\Big{\|}^{2}_{C_{n^{4}}\otimes_{min}\mathcal{B}({H})}$
$=\Big{\|}\sum_{\begin{subarray}{c}x,y,a,b\in\mathbb{Z}_{n}\\ a+b=xy\end{subarray}}T(e_{a,x})^{\dagger}T(e_{a,x})S(e_{b,y})^{\dagger}S(e_{b,y})\Big{\|}_{\mathcal{B}({H})},$
where we have used the commutativity relations between $T$ and $S$.
If we denote $E_{x}^{a}=T(e_{a,x})^{\dagger}T(e_{a,x})$, $F_{y}^{b}=S(e_{b,y})^{\dagger}S(e_{b,y})$ and $\omega=e^{\frac{2\pi i}{n}}$, we can write the previous expression as
$\Big{\|}\frac{1}{n}\sum_{k=0}^{n-1}\sum_{x,y,a,b\in\mathbb{Z}_{n}}\omega^{k(a+b-xy)}E_{x}^{a}F_{y}^{b}\Big{\|}_{\mathcal{B}({H})}=\frac{1}{n}\Big{\|}\sum_{k=0}^{n-1}\sum_{x,y\in\mathbb{Z}_{n}}\omega^{-kxy}\Big{(}\sum_{a}\omega^{ka}E_{x}^{a}\Big{)}\Big{(}\sum_{b}\omega^{kb}F_{y}^{b}\Big{)}\Big{\|}_{\mathcal{B}({H})}.$
According to Lemma 3.2, the complete contractivity of $T$ and $S$ guarantees that $T_{x}^{k}:=\sum_{a}\omega^{ka}E_{x}^{a}$ and $S_{y}^{k}:=\sum_{b}\omega^{kb}F_{y}^{b}$ are contractions in $\mathcal{B}({H})$ for every $x$, $y$, $k$. On the other hand, for every norm-one vectors $\xi,\,\eta\in H$, we have
$\frac{1}{n}\sum_{k=0}^{n-1}\sum_{x,y\in\mathbb{Z}_{n}}\omega^{-kxy}\langle\eta,(T_{x}^{k}S_{y}^{k})(\xi)\rangle=\frac{1}{n}\sum_{k=0}^{n-1}\langle v_{k},(\text{Id}\otimes\Phi_{k})(u_{k})\rangle,$
where $u_{k}=\sum_{y}S_{y}^{k}(\xi)\otimes e_{y}\in{H}\otimes\ell_{2}^{n}$, $v_{k}=\sum_{x}(T_{x}^{k})^{\dagger}(\eta)\otimes e_{x}\in{H}\otimes\ell_{2}^{n}$ and $\Phi_{k}=(\omega^{-kxy})_{x,y}\in M_{n}$.
Now, it is clear that $\|u_{k}\|\leq\sqrt{n}$ and $\|v_{k}\|\leq\sqrt{n}$ for every $k$. Also, $\|\Phi_{k}\|=\sqrt{n}$ for every $k\neq 0$ (note that $(1/\sqrt{n})\Phi_{k}$ is a unitary matrix) and $\|\Phi_{0}\|=n$. Hence, we obtain
$\frac{1}{n}\sum_{k=0}^{n-1}\langle v_{k},(\text{Id}\otimes\Phi_{k})(u_{k})\rangle\leq n+(n-1)\sqrt{n}\leq 2n\sqrt{n}.$
The following lemma will be very useful in order to prove the lower bound in Theorem 3.1.
Lemma 3.4.
For any natural number $n\geq 2$, let’s consider a function $f:\mathbb{Z}_{n}^{4}\rightarrow\{0,1\}$ such that for every $x$, $y$, there exists a bijection $\pi_{x,y}:\mathbb{Z}_{n}\rightarrow\mathbb{Z}_{n}$ such that $f(x,y,a,b)=1$ if and only if $\pi_{x,y}(a)=b$. Then,
$\xi_{f}=\sum_{x,y,a,b\in\mathbb{Z}_{n}}f(x,y,a,b)\,e_{xyab}\otimes e_{a,x}\otimes e_{b,y}$
satisfies
$\|\xi_{f}\|_{R_{n^{4}}\otimes_{h}(S_{\infty}^{n}\otimes_{h}S_{\infty}^{n})}={n^{2}}.$
The same estimate applies to the element
$\xi^{T}_{f}=\sum_{x,y,a,b\in\mathbb{Z}_{n}}f(x,y,a,b)\,e_{xyab}\otimes e_{b,y}\otimes e_{a,x}.$
Proof.
The homogeneity of the row operator spaces structure, together with the associativity of the Haagerup tensor norm and the completely isometric identification $S_{\infty}^{n}=C_{n}\otimes_{h}R_{n}$, allow us to rearrange the tensors $e_{xyab}=e_{x}\otimes e_{y}\otimes e_{a}\otimes e_{b}$ to write
$\|\xi_{f}\|_{R_{n^{4}}\otimes_{h}(S_{\infty}^{n}\otimes_{h}S_{\infty}^{n})}=\Big{\|}\sum_{y}e_{y}\otimes\xi_{y}\otimes e_{y}\Big{\|}_{R_{n}\otimes_{h}Y\otimes_{h}R_{n}},$
where $Y=R_{n^{3}}\otimes_{h}S_{\infty}^{n}\otimes_{h}C_{n}$ and
$\xi_{y}=\sum_{x,a,b}f(x,y,a,b)e_{xab}\otimes e_{a,x}\otimes e_{b}\in Y.$
According to Lemma 2.3, we have
$\|\xi_{f}\|_{R_{n^{4}}\otimes_{h}(S_{\infty}^{n}\otimes_{h}S_{\infty}^{n})}=\Big{(}\sum_{y}\|\xi_{y}\|_{Y}^{2}\Big{)}^{\frac{1}{2}}.$
In addition, for any fixed $y$ we can again use the properties of the Haagerup tensor norm to write
$\|\xi_{y}\|_{Y}=\Big{\|}\sum_{b}e_{b}\otimes\xi_{y,b}\otimes e_{b}\Big{\|}_{R_{n}\otimes_{h}Z\otimes_{h}C_{n}},$
where $Z=R_{{n^{2}}}\otimes_{h}S_{\infty}^{n}$ and
$\xi_{y,b}=\sum_{x,a}f(x,y,a,b)e_{xa}\otimes e_{a,x}\in Z.$
Now, according to Lemma 2.4, we have
$\|\xi_{y}\|_{Y}=\sum_{b}\Big{\|}\sum_{x,a}f(x,y,a,b)e_{xa}\otimes e_{a,x}\Big{\|}_{Z}.$
Finally, by once again invoking Lemma 2.3, we find that for ever $y$ and $b$, we have
$\Big{\|}\sum_{x,a}f(x,y,a,b)e_{xa}\otimes e_{a,x}\Big{\|}_{Z}$ $=\Big{\|}\sum_{x}e_{x}\otimes\Big{(}\sum_{a}f(x,y,a,b)e_{a}\otimes e_{a}\Big{)}\otimes e_{x}\Big{\|}_{R_{n}\otimes_{h}(S_{1}^{n})\otimes_{h}R_{n}}$
$=\Big{(}\sum_{x}\Big{\|}\sum_{a}f(x,y,a,b)e_{a}\otimes e_{a}\Big{\|}_{S_{1}^{n}}^{2}\Big{)}^{\frac{1}{2}}=\sqrt{n},$
where the last equality follows from the properties of the function $f$.
It then follows that $\|\xi_{f}\|_{R_{n^{4}}\otimes_{h}(S_{\infty}^{n}\otimes_{h}S_{\infty}^{n})}=n^{2}$ as we wanted.
The estimate for $\xi^{T}_{f}$ can be proved completely analogously by just interchanging $x$,$y$ and $a$,$b$. ∎
We are now ready to prove Theorem 3.1. To do so, we will use notation ${H}_{A}={H}_{B}=\ell_{2}^{n}$ to indicate the order considered when taking the Haagerup tensor norm.
Proof of Theorem 3.1.
According to Lemma 3.3, for any prime number $n$ the element
$\eta_{n}=\sum_{\begin{subarray}{c}x,y,a,b\in\mathbb{Z}_{n}\\ a+b=xy\end{subarray}}e_{xyab,1}\otimes e_{a,x}\otimes e_{b,y}$
satisfies
$\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}(H_{A})\otimes_{max}S_{1}(H_{B}))}\leq\sqrt{2}n^{\frac{3}{4}}.$
On the other hand, according to Lemma 3.4 applied to the function $f(x,y,a,b)=\delta_{b,xy-a}$, the element
$P_{n}=\frac{1}{{n^{2}}}\sum_{\begin{subarray}{c}x,y,a,b\in\mathbb{Z}_{n}\\ a+b=xy\end{subarray}}e_{1,xyab}\otimes e_{a,x}\otimes e_{b,y}$
satisfies
(3.2) $\max\{\|P_{n}\|_{R_{n^{4}}\widehat{\otimes}(S_{\infty}(H_{A})\otimes_{h}S_{\infty}(H_{B}))},\|P_{n}^{T}\|_{R_{n^{4}}\widehat{\otimes}(S_{\infty}(H_{B})\otimes_{h}S_{\infty}(H_{A}))}\}=1.$
Hence, for any decomposition $\eta_{n}=\eta_{n,1}+\eta_{n,2}$ in $C_{n^{4}}\otimes S_{1}(H_{A})\otimes S_{1}(H_{B})$ we have that
$n$ $=\langle\eta_{n},P_{n}\rangle=\langle\eta_{n,1},P_{n}\rangle+\langle\eta_{n,2},P_{n}\rangle=\langle\eta_{n,1},P_{n}\rangle+\langle\eta_{n,2}^{T},P_{n}^{T}\rangle$
$\leq\|\eta_{n,1}\|_{C_{n^{4}}\otimes_{min}(S_{1}(H_{A})\otimes_{h}S_{1}(H_{B}))}+\|\eta_{n,2}^{T}\|_{C_{n^{4}}\otimes_{min}(S_{1}(H_{B})\otimes_{h}S_{1}(H_{A}))}.$
According to the definition of the $\mu$ norm (2.4), this concludes the proof. ∎
4. Proof of Theorem 1.2
In this section we prove the following theorem, from which Theorem 1.2 follows immediately.
Theorem 4.1.
For every prime number $n$ there exists an element $\beta_{n}\in M_{n^{6}}\otimes\ell_{1}^{n^{2}}\otimes\ell_{1}^{n^{2}}$ such that $\|\eta_{n}\|_{S_{\infty}^{n^{6}}\otimes_{min}(\ell_{1}^{n^{2}}\otimes_{max}\ell_{1}^{n^{2}})}\leq\sqrt{2}n^{\frac{3}{4}}$ and $\|\eta_{n}\|_{S_{\infty}^{n^{6}}\otimes_{min}(\ell_{1}^{n^{2}}\otimes_{\mu}\ell_{1}^{n^{2}})}\geq n$.
The strategy to prove Theorem 1.2 consists of transforming the element $\eta_{n}\in C\otimes S_{1}\otimes S_{1}$ defined in the proof of Theorem 1.1, into another element $\beta_{n}\in S_{\infty}\otimes\ell_{1}\otimes\ell_{1}$ such that the norms $max$ and $\mu$ are preserved. To do this, given $n$, let us consider the unitary matrices $X,\,Z\in M_{n}$, defined by $X(e_{j})=e^{\frac{2\pi ij}{n}}e_{j}$ and $Z(e_{j})=e_{j+1}$ respectively on the canonical vectors $(e_{j})_{j=0}^{n-1}$ and denote $T_{k,l}=X^{k}Z^{l}$ for every $k,l=0,1,\ldots,n-1$.
The following lemma, whose proof is included for completeness, was proved in [12] and is inspired by the quantum teleportation protocol in quantum information theory.
Lemma 4.2.
Let $J:S_{1}^{n}\rightarrow S_{\infty}^{n}(\ell_{1}^{n^{2}})$ and $W:S_{\infty}^{n}\rightarrow S_{\infty}^{n}(\ell_{\infty}^{n^{2}})$ be the linear maps defined, respectively, as
$J(\rho)=\frac{1}{n}\sum_{k,l=0}^{n-1}T_{k,l}^{\dagger}\rho T_{k,l}\otimes e_{k,l}\hskip 8.5359pt\text{and}\hskip 8.5359ptW(\rho)=\sum_{k,l=0}^{n-1}T_{k,l}^{T}\rho\overline{T}_{k,l}\otimes e_{k,l}.$
Then, $J$ and $W$ are complete contractions.
Proof.
Since $W$ is a completely positive and unital map between $C^{*}$-algebras, we have $\|W\|_{cb}=1$.
In order to study the map $J$, let us define $\eta_{k,l}=(T_{k,l}\otimes\text{Id})(\phi)\in\ell_{2}^{{n^{2}}}$ for every $k,l=0,\ldots,n-1$, where $\phi=(1/\sqrt{n})\sum_{j=0}^{n-1}e_{j}\otimes e_{j}$. It is very easy to check that these vectors form an orthonormal basis of $\ell_{2}^{{n^{2}}}$. Moreover, one can check that for every $h\in\ell_{2}^{n}$, the following identity holds
$h\otimes\sum_{i=0}^{n-1}e_{i}\otimes e_{i}=\frac{1}{\sqrt{n}}\sum_{k,l=0}^{n-1}\eta_{k,l}\otimes T_{k,l}^{\dagger}(h).$
This allows us to write, for every $\rho\in S_{1}^{n}$, the identity
$\rho\otimes\phi=\frac{1}{n}\sum_{k,l,k^{\prime},l^{\prime}=0}^{n-1}H_{k,l}^{k^{\prime},l^{\prime}}\otimes T_{k,l}^{\dagger}\rho T_{k^{\prime},l^{\prime}},$
where $\phi=\sum_{i,j=0}^{n-1}e_{i,j}\otimes e_{i,j}$ and $H_{k,l}^{k^{\prime},l^{\prime}}:M_{{n^{2}}}\rightarrow M_{{n^{2}}}$ is the rank one operator defined as $H_{k,l}^{k^{\prime},l^{\prime}}(v)=\langle\eta_{k^{\prime},l^{\prime}},v\rangle\eta_{k,l}$.
On the other hand, the fact that the vectors $\eta_{k,l}$, $k,l=0,1,\ldots,n-1$ form an orthonormal basis of $\ell_{2}^{{n^{2}}}$ guarantees that the projection $P:S_{1}^{{n^{2}}}\rightarrow\ell_{1}^{n^{2}}$, defined by
$P(A)=\sum_{k,l=0}^{n-1}\langle\eta_{k,l},A(\eta_{k,l})\rangle e_{k,l},$
is a complete contraction. According to Lemma 2.1, the identity map $\text{Id}:S_{1}^{n}\hat{\otimes}(S_{1}^{n}\otimes_{min}S_{\infty}^{n})\rightarrow S_{1}^{{n^{2}}}\otimes_{min}S_{\infty}^{n}$ is completely contractive. It then follows that the map $i:S_{1}^{n}\rightarrow S_{1}^{{n^{2}}}\otimes_{min}S_{\infty}^{n}$, defined by $i(\rho)=\rho\otimes\phi$, is automatically completely contractive and the complete contractivity of $J$ follows from the identity $J=(P\otimes\text{Id})\circ i$. ∎
Lemma 4.2 is the key point to transfer the statement of Theorem 3.1 on non-commutative $L_{1}$-spaces into a statement involving commutative ones. Indeed, if $\eta_{n}$ is the element defined in (3.1), then we define $\beta_{n}=(\text{Id}\otimes J\otimes J)(\eta_{n})$, so that
(4.1) $\beta_{n}=\frac{1}{n^{2}}\sum_{\begin{subarray}{c}x,y,a,b\in\mathbb{Z}_{n}\\ a+b=xy\\ k,l,k^{\prime},l^{\prime}\in\mathbb{Z}_{n}\end{subarray}}e_{xyab,1}\otimes R_{a,x}^{k,l}\otimes R_{b,y}^{k^{\prime},l^{\prime}}\otimes e_{k,l}\otimes e_{k^{\prime},l^{\prime}}\in C_{n^{4}}\otimes S_{\infty}^{n}\otimes S_{\infty}^{n}\otimes(\ell_{1}^{n^{2}}\otimes\ell_{1}^{n^{2}}),$
where we have denoted $R_{a,x}^{k,l}=T_{k,l}^{\dagger}e_{a,x}T_{k,l}$ for every $a,x,k,l$.
Note that the complete contractivity of $J$ along with the commutative and associative properties of the min norm and the metric mapping property immediately imply that
$\|\beta_{n}\|_{C_{n^{4}}\otimes_{min}S_{\infty}^{n^{2}}\otimes_{min}(\ell_{1}^{n^{2}}\otimes_{min}\ell_{1}^{n^{2}})}\leq\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{min}S_{1}^{n})}.$
The next lemma shows that we can actually upper bound the max norm of $\beta_{n}$.
Lemma 4.3.
Given any natural number $n$, we have
$\|\beta_{n}\|_{C_{n^{4}}\otimes_{min}S_{\infty}^{n^{2}}\otimes_{min}(\ell_{1}^{n^{2}}\otimes_{max}\ell_{1}^{n^{2}})}\leq\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{max}S_{1}^{n})}.$
Proof.
Given any pair of (complete) contractions $T,S:\ell_{1}^{n^{2}}\rightarrow\mathcal{B}({H})$ such that $[T(e_{k,l}),S(e_{k^{\prime},l^{\prime}})]=[T(e_{k,l}),S(e_{k^{\prime},l^{\prime}})^{\dagger}]=0$ for every $k,l,k^{\prime},l^{\prime}$, we define the linear maps
$\tilde{T},\tilde{S}:M_{n}(\ell_{1}^{n^{2}})\rightarrow M_{n}\otimes_{min}M_{n}\otimes_{min}\mathcal{B}({H})=\mathcal{B}(\ell_{2}^{n}\otimes_{2}\ell_{2}^{n}\otimes_{2}{H}),$
as $\tilde{T}\Big{(}\sum_{k,l}A_{k,l}\otimes e_{k,l}\Big{)}=\sum_{k,l}A_{k,l}\otimes 1\!\!1\otimes T(e_{k,l})$ and $\tilde{S}\Big{(}\sum_{k,l}B_{k,l}\otimes e_{k,l}\Big{)}=\sum_{k,l}1\!\!1\otimes B_{k,l}\otimes S(e_{k,l})$.
Since $\tilde{T}$ and $\tilde{S}$ are clearly complete contractions, we deduce from Lemma 4.2 that the maps
$\hat{T},\hat{S}:S_{1}^{n}\rightarrow\mathcal{B}(\ell_{2}^{n}\otimes_{2}\ell_{2}^{n}\otimes_{2}{H}),$
defined as $\hat{T}=\tilde{T}\otimes J$ and $\hat{S}=\tilde{S}\otimes J$ respectively, are also completely contractive. Moreover, it is clear that $[\hat{T}(\rho),\hat{S}(\gamma)]=[\hat{T}(\rho),\hat{S}(\gamma)^{{\dagger}}]=0$ for every $\rho,\,\gamma\in S_{1}^{n}$. Then, we conclude our proof by noticing that
$\|(\text{Id}_{C_{n^{4}}\otimes S_{\infty}^{n^{2}}}\otimes T\odot S)(\beta_{n})\|_{C_{n^{4}}\otimes_{min}S_{\infty}^{n}\otimes_{min}S_{\infty}^{n}\otimes_{min}\mathcal{B}({H})}$
$=\|(\text{Id}_{C_{n^{4}}}\otimes\hat{T}\odot\hat{S})(\eta_{n})\|_{C_{n^{4}}\otimes_{min}\mathcal{B}(\ell_{2}^{n}\otimes_{2}\ell_{2}^{n}\otimes_{2}{H})}$
$\leq\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}^{n}\otimes_{max}S_{1}^{n})}.$
We are now ready to prove the main result of the paper. It’s very useful to denote ${H}_{A}={H}_{B}=\ell_{2}^{n}$ as well as $\ell_{\infty}^{{n_{A}}^{2}}=\ell_{\infty}^{{n_{B}}^{2}}=\ell_{\infty}^{{n}^{2}}$ to precisely indicate the rearrangement in the various steps of the proof.
Proof of Theorem 1.2.
On the one hand, according to Lemma 4.3, we have that
$\|\beta_{n}\|_{C_{n^{4}}\otimes_{min}S_{\infty}(H_{A})\otimes_{min}S_{\infty}(H_{B})\otimes_{min}(\ell_{1}^{n_{A}^{2}}\otimes_{max}\ell_{1}^{n_{B}^{2}})}\leq\|\eta_{n}\|_{C_{n^{4}}\otimes_{min}(S_{1}(H_{A})\otimes_{max}S_{1}(H_{B}))}\leq\sqrt{2}n^{\frac{3}{4}}.$
On the other hand, using the notation of Lemma 4.2, let us consider the element
(4.2) $Q_{n}=\frac{1}{n^{2}}(\text{Id}_{R_{n^{4}}}\otimes W\otimes W)(P_{n})=\frac{1}{n^{4}}\sum_{\begin{subarray}{c}x,y,a,b\in\mathbb{Z}_{n}\\ a+b=xy\\ k,l,k^{\prime},l^{\prime}\in\mathbb{Z}_{n}\end{subarray}}e_{1,xyab}\otimes S_{a,x}^{k,l}\otimes S_{b,y}^{k^{\prime},l^{\prime}}\otimes e_{k,l}\otimes e_{k^{\prime},l^{\prime}},$
in $R_{n^{4}}\otimes S_{\infty}(H_{A})\otimes S_{\infty}(H_{B})\otimes(\ell_{\infty}^{n_{A}^{2}}\otimes\ell_{\infty}^{n_{B}^{2}})$, where we have denoted $S_{a,x}^{k,l}=T_{k,l}^{T}e_{a,x}\overline{T}_{k,l}$ for every $a,x,k,l$.
Since $\langle\beta_{n},Q_{n}\rangle=\langle\eta_{n},P_{n}\rangle=n$, we can conclude the proof of the theorem in the same way as the proof of Theorem 1.1, by simply demonstrating that
$\max\left\{\|Q_{n}\|_{(R_{n^{4}}\widehat{\otimes}S_{1}(H_{A})\widehat{\otimes}S_{1}(H_{B}))\widehat{\otimes}(\ell_{\infty}^{{n_{A}}^{2}}\otimes_{h}\ell_{\infty}^{{n_{B}}^{2}})},\|Q_{n}^{T}\|_{(R_{n^{4}}\widehat{\otimes}S_{1}(H_{A})\widehat{\otimes}S_{1}(H_{B}))\widehat{\otimes}(\ell_{\infty}^{{n_{B}}^{2}}\otimes_{h}\ell_{\infty}^{{n_{A}}^{2}})}\right\}\leq 1.$
In order to prove the inequality
$\|Q_{n}\|_{(R_{n^{4}}\widehat{\otimes}S_{1}(H_{A})\widehat{\otimes}S_{1}(H_{B}))\widehat{\otimes}(\ell_{\infty}^{{n_{A}}^{2}}\otimes_{h}\ell_{\infty}^{{n_{B}}^{2}})}\leq 1,$
we used that $\|P_{n}\|_{R_{n^{4}}\widehat{\otimes}(S_{\infty}(H_{A})\otimes_{h}S_{\infty}(H_{B}))}\leq 1$ and the fact that $W$ is a complete contraction to guarantee that
$\|(\text{Id}_{R_{n^{4}}}\otimes W\otimes W)(P_{n})\|_{R_{n^{4}}\widehat{\otimes}\big{[}(S_{\infty}(H_{A})\otimes_{min}\ell_{\infty}^{{n_{A}}^{2}})\otimes_{h}(S_{\infty}(H_{B})\otimes_{min}\ell_{\infty}^{{n_{B}}^{2}})\big{]}}\leq 1.$
Now, according to the first property proved in Lemma 2.2 applied to $X=S_{\infty}(H_{A})\otimes_{min}\ell_{\infty}^{{n_{A}}^{2}}$ and $Y=\ell_{\infty}^{{n_{B}}^{2}}$, the previous estimate implies that
$\|(\text{Id}_{R_{n^{4}}}\otimes W\otimes W)(P_{n})\|_{R_{n^{4}}\widehat{\otimes}\Big{[}S_{\infty}(H_{B})\otimes_{min}\big{(}\big{(}S_{\infty}(H_{A})\otimes_{min}\ell_{\infty}^{{n_{A}}^{2}}\big{)}\otimes_{h}\ell_{\infty}^{{n_{B}}^{2}}\big{)}\Big{]}}\leq 1,$
where the terms must be rearranged according to the map $T$ in Lemma 2.2.
We can now apply the second property proven in Lemma 2.2 to $X=\ell_{\infty}^{{n_{A}}^{2}}$ and $Y=\ell_{\infty}^{{n_{B}}^{2}}$ to conclude that the previous estimate implies that
$\|(\text{Id}_{R_{n^{4}}}\otimes W\otimes W)(P_{n})\|_{R_{n^{4}}\widehat{\otimes}\Big{[}S_{\infty}(H_{B})\otimes_{min}S_{\infty}(H_{A})\otimes_{min}\big{(}\ell_{\infty}^{{n_{A}}^{2}}\otimes_{h}\ell_{\infty}^{{n_{B}}^{2}}\big{)}\Big{]}}\leq 1.$
Finally, by using that $S_{\infty}(H_{A})=S_{\infty}(H_{B})=S_{\infty}^{n}$ and Equation (2.3), we can conclude that
$\|Q_{n}\|_{R_{n^{4}}\widehat{\otimes}S_{1}(H_{B})\widehat{\otimes}S_{1}(H_{A})\widehat{\otimes}\Big{(}\ell_{\infty}^{{n_{A}}^{2}}\otimes_{h}\ell_{\infty}^{{n_{B}}^{2}}\Big{)}}\leq 1.$
Note that the commutativity of the projective tensor norms implies that
$\|Q_{n}\|_{R_{n^{4}}\widehat{\otimes}S_{1}(H_{A})\widehat{\otimes}S_{1}(H_{B})\widehat{\otimes}\Big{(}\ell_{\infty}^{{n_{A}}^{2}}\otimes_{h}\ell_{\infty}^{{n_{B}}^{2}}\Big{)}}\leq 1$
as we wanted.
Given that $\|P_{n}^{T}\|_{R_{n^{4}}\widehat{\otimes}(S_{\infty}(H_{B})\otimes_{h}S_{\infty}(H_{A}))}\leq 1$, employing the same argument enables us to deduce that
$\|Q_{n}^{T}\|_{R_{n^{4}}\widehat{\otimes}S_{1}(H_{A})\widehat{\otimes}S_{1}(H_{B})\widehat{\otimes}\Big{(}\ell_{\infty}^{{n_{B}}^{2}}\otimes_{h}\ell_{\infty}^{{n_{A}}^{2}}\Big{)}}\leq 1.$
This concludes the proof. ∎
Remark 4.1.
Let $(x_{i})_{i}$ and $(y_{j})_{j}$ be two families of contractions in $\mathcal{B}(H)$ such that $[x_{i},y_{j}]=0$ for every $i,j$. If we assume, in addition, the elements $x_{j}$ to be normal, it follows from Fuglede’s Theorem [7] that $[x_{i},y_{j}^{\dagger}]=0$ for every $i,j$. According to Russo-Dye Theorem, we can then write that for every family of matrices $(a_{i,j})_{i,j}\in M_{n}$, it holds that
$\Big{\|}\sum_{i,j}a_{i,j}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\,\leq\,\sup\Big{\|}\sum_{i,j}a_{ij}\otimes u_{i}v_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\leq\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
where the supremum runs over all families of commuting unitaries $(u_{i})_{i}$ ,$(v_{j})_{j}$ in $\mathcal{B}(H)$. While Ricard demonstrated [19, Proposition 13.10] the existence of families of commuting contractions $(x_{i})_{i}$ ,$(y_{j})_{j}$ satisfying
$\Big{\|}\sum_{i,j}a_{i,j}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}>\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
even for scalar coefficients $(a_{i,j})_{i,j}\in\mathbb{C}$ for every $i,j$, Theorem 1.2 provides families of commuting contractions that dramatically violate this inequality - such that even an equivalence between the norms is not possible. In this sense, Theorem 1.2 can be understood as a qualitative witness to the fact that these commuting contractions cannot be normal. We are grateful to Mikael $\o$rdam for reminding us of Fuglede’s result.

### v2

@@
 (1.1) $\ell_{1}\otimes_{\epsilon}\ell_{1}\cong\ell_{1}\otimes_{\gamma_{2}^{*}}\ell_{1}\quad\mbox{isomorphically}\,,$
-where $\gamma_{2}^{*}$ is dual of the $\gamma_{2}$ norm, derived from the Hilbert space factorization norm for linear maps.
+where $\gamma_{2}^{*}$ is dual norm of the Hilbert space factorization norm $\gamma_{2}$.
 This theorem highlights a remarkable property of tensor products of Banach spaces, revealing deep connections between different norms and structures (see [5] and the references therein).
-In 1974 Pisier provided a positive answer to one of Grothendieck’s problems by showing that
-(1.2) $A^{*}\otimes_{\epsilon}B^{*}\cong A^{*}\otimes_{\gamma_{2}^{*}}B^{*}\quad\mbox{isomorphically}$
-holds for $C^{*}$-algebra $A$, $B$ with an additional approximability assumption [16].
+In particular, Grothendieck’s theorem implies that every linear map $T:\ell_{\infty}\rightarrow\ell_{1}$ factors through a Hilbert space.
+In 1974, Pisier [16] provided a positive answer to one of Grothendieck’s conjectures by showing that
+(1.2) $A^{*}\otimes_{\epsilon}B^{*}\cong A^{*}\otimes_{\tilde{\gamma}^{*}_{2}}B^{*}\quad\mbox{isomorphically}$
+holds for general $C^{*}$-algebras $A$ and $B$.
+Here, $\tilde{\gamma}^{*}_{2}$ is the analog to the $\gamma_{2}^{*}$ norm in the context of operator algebras.
 That is, Pisier replaced commutative $L_{1}$-spaces by noncommutative ones.
-This result was later improved by Haagerup [9], who removed the approximability assumption made in [16] and clarified the correct analogue to $\gamma_{2}^{*}$ in the context of operator algebras.
+In particular, Pisier proved that every linear map $T:A\rightarrow B^{*}$ which is approximable, in a particular way, by finite ranks maps, factors through a Hilbert space.
+This result was later extended by Haagerup [9], who removed such an approximability assumption.
 The category of operator spaces, which are closed subspaces of $C^{*}$-algebras endowed with an induced matricial structure, is closed under taking dual spaces, closed subspaces and quotients.
@@
 (1.3) $A^{*}\otimes_{\min}B^{*}\,\stackrel{{\scriptstyle?}}{{\cong}}A^{*}\otimes_{\mu}B^{*}\quad\mbox{completely isomorphically}\,,$
-where $A^{*}\otimes_{\mu}B^{*}=A^{*}\otimes_{h}B^{*}+B^{*}\otimes_{h}A^{*}$ may be considered as a quotient of the direct sum $A^{*}\otimes_{h}B^{*}\oplus B^{*}\otimes_{h}A^{*}$ with respect to the map $q(a_{1}\otimes b_{1}\oplus b_{2}\otimes a_{2})=a_{1}\otimes b_{1}+a_{2}\otimes b_{2}$.
-A breakthrough in this problem was achieved by Piser and Shlyakhtenko in [20], where they conclusively demonstrated that the answer to the preceding question is indeed positive at the Banach space level.
+where $A^{*}\otimes_{\mu}B^{*}=A^{*}\otimes_{h}B^{*}+B^{*}\otimes_{h}A^{*}$ may be considered as a quotient of the direct sum $A^{*}\otimes_{h}B^{*}\oplus_{1}B^{*}\otimes_{h}A^{*}$ with respect to the map $q(a_{1}\otimes b_{1}\oplus b_{2}\otimes a_{2})=a_{1}\otimes b_{1}+a_{2}\otimes b_{2}$.
+A breakthrough in this problem was achieved by Pisier and Shlyakhtenko in [20], where they conclusively demonstrated that the answer to the preceding question is indeed positive at the Banach space level.
 They proved
 (1.4) $A^{*}\otimes_{\min}B^{*}\cong A^{*}\otimes_{\mu}B^{*}\quad\mbox{isomorphically}\,,$
-assuming that the $C^{*}$-algebras $A$ and $B$ are exact.
+for general $C^{*}$-algebras $A$ and $B$.
 Notably, their results extended beyond the setting of $C^{*}$-algebras to encompass general exact operator spaces.
-Subsequently, Haagerup and Musat in [10] furthered this understanding to include general $C^{*}$-algebras.
-However, the broader question concerning the operator space structure of the corresponding tensors products in Conjecture (1.3) remained unresolved in [20] (see also [18, Problem 21.2]).
+Subsequently, Haagerup and Musat in [10] furthered the understanding of (1.4), particularly improving the constant involved.
+However, the broader question concerning the operator space structure of the corresponding tensor products in Conjecture (1.3) remained unresolved in [20] (see also [18, Problem 21.2]).
 The first result presented in this work provides a negative answer to Blecher’s conjecture.
@@
 Theorem 1.1.
-$C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\min}\mathcal{K}(\ell_{2})^{*})\neq C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\mu}\mathcal{K}(\ell_{2})^{*})\quad\text{isomorphically}\,,$
+$C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\min}\mathcal{K}(\ell_{2})^{*})\neq C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\mu}\mathcal{K}(\ell_{2})^{*})\quad\text{isomorphically}\,.$
 In fact, we will demonstrate an even stronger result than Theorem 1.1 by proving the same statement with the operator space $\mathcal{K}(\ell_{2})^{*}\otimes_{\min}\mathcal{K}(\ell_{2})^{*}$ replaced by $\mathcal{K}(\ell_{2})^{*}\otimes_{\max}\mathcal{K}(\ell_{2})^{*}$.
@@
 Since it is well known that $C^{*}\langle\ell_{1}\rangle=C^{*}(\mathbb{F}_{\infty})$, one can alternatively express Theorem 1.2 by indicating that the natural inclusion $OA(\ell_{1})\otimes_{max}OA(\ell_{1})\subseteq C^{*}(\mathbb{F}_{\infty})\otimes_{max}C^{*}(\mathbb{F}_{\infty})$ is not completely isomorphic, even when restricted to the linear subspace $\ell_{1}\otimes\ell_{1}$.
-According to the precise definitions of $OA(\ell_{1})$ and $C^{*}\langle\ell_{1}\rangle$, Theorem 1.2 underscores a significant limitation in lifting specific sets of commuting contractions to commuting unitaries (see Remark 4.1).
+According to the precise definitions of $OA(\ell_{1})$ and $C^{*}\langle\ell_{1}\rangle$, Theorem 1.2 underscores a significant limitation involving commuting contractions and commuting unitaries, or even commuting normal contractions (see Remark 4.1).
 The approach followed in this work delves into the deep relationship between the theories of operator spaces and nonlocal games [15].
@@
 Let $(x_{i})_{i}$ and $(y_{j})_{j}$ be two families of contractions in $\mathcal{B}(H)$ such that $[x_{i},y_{j}]=0$ for every $i,j$.
-If we assume, in addition, the elements $x_{j}$ to be normal, it follows from Fuglede’s Theorem [7] that $[x_{i},y_{j}^{\dagger}]=0$ for every $i,j$.
-According to Russo-Dye Theorem, we can then write that for every family of matrices $(a_{i,j})_{i,j}\in M_{n}$, it holds that
-$\Big{\|}\sum_{i,j}a_{i,j}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\,\leq\,\sup\Big{\|}\sum_{i,j}a_{ij}\otimes u_{i}v_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\leq\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
+If we assume, in addition, the elements $x_{i}$’s to be normal, it follows from Fuglede’s Theorem [7] that $[x_{i},y_{j}^{\dagger}]=0$ for every $i,j$.
+According to Russo-Dye Theorem (or by just using twice the standard block encoding; that is, given any contraction $T\in\mathcal{B}(H)$, we consider the unitary
+$U\,=\,\left(\begin{array}[]{cc}T&-\sqrt{1-TT^{\dagger}}\\ \sqrt{1-T^{\dagger}T}&T^{\dagger}\end{array}\right)\in M_{2}(\mathcal{B}(H))),$
+we obtain that for every family of matrices $(a_{ij})_{i,j}\in M_{n}$, it holds that
+$\Big{\|}\sum_{i,j}a_{ij}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\,\leq\,\sup\Big{\|}\sum_{i,j}a_{ij}\otimes u_{i}v_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\leq\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
 where the supremum runs over all families of commuting unitaries $(u_{i})_{i}$, $(v_{j})_{j}$ in $\mathcal{B}(H)$.
-While Ricard demonstrated [19, Proposition 13.10] the existence of families of commuting contractions $(x_{i})_{i}$ ,$(y_{j})_{j}$ satisfying
-$\Big{\|}\sum_{i,j}a_{i,j}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}>\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
-even for scalar coefficients $(a_{i,j})_{i,j}\in\mathbb{C}$ for every $i,j$, Theorem 1.2 provides families of commuting contractions that dramatically violate this inequality - such that even an equivalence between the norms is not possible.
-In this sense, Theorem 1.2 can be understood as a qualitative witness to the fact that these commuting contractions cannot be normal.
+While Ricard demonstrated [19, Proposition 13.10] the existence of families of commuting contractions $(x_{i})_{i}$, $(y_{j})_{j}$ and (even scalar) coefficients $(a_{ij})_{i,j}\in\mathbb{C}$ satisfying
+$\Big{\|}\sum_{i,j}a_{ij}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}>\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
+Theorem 1.2 provides families of commuting contractions that dramatically violate this inequality - such that even an equivalence between the norms is not possible.
+Note that considering matrix coefficients $(a_{ij})_{i,j}$ is crucial here, since Grothendieck’s inequality (1.1) implies that
+$\Big{\|}\sum_{i,j}a_{ij}x_{i}y_{j}\Big{\|}_{\mathcal{B}(H)}\leq K_{G}\Big{\|}\sum_{ij}a_{ij}e_{i}\otimes e_{j}\Big{\|}_{\ell_{1}\otimes_{\max}\ell_{1}},\,$
+for every families of (non-necessarily commuting) contractions $(x_{i})_{i}$ and $(y_{j})_{j}$.
+Hence, Theorem 1.2 can be understood as a qualitative witness to the fact that the commuting contractions $(x_{i})_{i}$ and $(y_{j})_{j}$ cannot be normal.
 We are grateful to Mikael $\o$rdam for reminding us of Fuglede’s result.